OpenStudy (anonymous):

Solve the equation. x2 = 96

Nnesha (nnesha):

2 is exponent

Nnesha (nnesha):

\[x^2 = 96\]

Nnesha (nnesha):

again @Soniatonia do you know difference of square method ??

OpenStudy (ilovecake):

http://www.tiger-algebra.com/drill/x2=96/

OpenStudy (anonymous):

not really

Nnesha (nnesha):

did you get the answer for last question ???

OpenStudy (anonymous):

yeah

Nnesha (nnesha):

well i'm wonder if you don't know difference of square method how did you get answer there i'll go with easy one take square root both side and let me know what you get

Nnesha (nnesha):

got it ??

OpenStudy (anonymous):

yes

Nnesha (nnesha):

what you got ??

Nnesha (nnesha):

@sammixboo

sammixboo (sammixboo):

Well, \(x^2 = 96\) is the same as \(x = \sqrt{\pm96}\)

sammixboo (sammixboo):

omg put the positive/negative sign in the wrong place

sammixboo (sammixboo):

Well, \(x^2 = 96\) is the same as \(x = \pm\sqrt{96}\)******

OpenStudy (anonymous):

ok